Let B = {(1,0), (0, 1)} and B = {(0, 1), (1, 1)} be two bases for the vector space V = RP. Moreover, let [y]g = [1 -2] and
Answer #1

B={(1,0), (011)}, and B = {(0,1), (1; 1)} {(0,1), (1, 1)} be two basis for v=1R² [-] the matrix relative to B is (TJO et and5 [w] B PB7B [u] B [!]-[ 2 - Bfrom [A], we have T(1,0) = 2(1,0)+2(0,1)=(2,-2) +(0,1) = 211,0) + 2(0,1)= (2,2) T(x,y) = XT(1,0)+ y T(011) = Traiy) = (2x+2y,